# Edge-Cloud Collaborative Computing on Distributed Intelligence and Model Optimization: A Survey

### v4

@@ III-A2 Quantization
Precision reduction techniques effectively minimize memory footprint and computational cost by employing lower-precision representations of model parameters [74], making quantized models particularly suitable for resource-constrained devices. Instead of 32-bit floating-point representations, quantization employs lower precision formats such as 8-bit integers, consequently diminishing memory requirements and accelerating computations on these lower-precision numbers. For example, [73] explored model quantization for edge implementation in arrhythmia monitoring to reduce energy consumption. Additionally, QuAsyncFL [68] integrates quantization with asynchronous federated learning to enhance communication efficiency in cloud-edge-terminal collaborations within the AIoT domain.
@@ III-A3 Knowledge Distillation
Knowledge transfer approaches enable compact ”student” models to achieve comparable performance with reduced complexity by learning from larger ”teacher” models [78], facilitating efficient deployment in resource-constrained environments [65]. Instead of learning from ground truth labels, the student model learns the teacher model’s output distribution, providing richer information than hard labels and enabling more nuanced representations and better generalization [69]. KD benefits ECCC by placing the teacher model in the resource-rich cloud and the distilled student model on edge devices [6], optimizing overall system performance. Agglomerative federated learning (FedAgg) [77] uses a bridge sample based online distillation protocol (BSBODP) for knowledge transfer between nodes across different tiers, enabling larger model training while maintaining privacy and flexibility. Guidance-based knowledge transfer (GKT) [78] employs a larger LLM as a ”teacher” to generate guidance prompts for a smaller ”student” model, enhancing accuracy and speed without fine-tuning. Growth-adaptive distillation [6] automatically adjusts the student model structure for efficient, low-power edge deployment. In cloud-edge collaborative reinforcement learning, algorithm-independent knowledge distillation [79] accelerates online RL agent convergence by leveraging pre-trained cloud models and offline agent interactions. A federated bi-directional knowledge distillation strategy [80] enhances fault detection in manufacturing by integrating quality knowledge from both cloud and edge, strengthening interaction between sub-processes and hierarchical levels.
@@ III-B Model Adaptation
Adaptive learning paradigms enable AI models to effectively respond to new tasks, datasets, or evolving data distributions within dynamic distributed computing environments. To achieve this adaptability, three primary learning paradigms are widely employed: transfer learning (TL), which accelerates learning on target tasks by leveraging knowledge from pre-trained models; federated learning (FL), which facilitates privacy-preserving, distributed training across multiple clients; and continual learning (CL), which enables models to learn sequentially from evolving data streams without catastrophic forgetting. The interplay and distinct roles of these paradigms in fostering adaptability within ECCC systems are conceptually depicted in Fig. 4.
@@
While some taxonomies treat federated learning as a distinct resource management or privacy technique, our classification positions FL within model adaptation because it fundamentally addresses how models adapt to distributed data patterns and heterogeneous environments, which is a core adaptation challenge in ECCC systems. Specifically: (1) Domain adaptation techniques (transfer learning) enable models to adapt across different domains or tasks by leveraging pre-existing knowledge; (2) Distributed adaptation methods (federated learning) enable models to adapt collaboratively across multiple data sources while preserving privacy and handling non-IID distributions; (3) Temporal adaptation approaches (continual learning) enable models to adapt incrementally to evolving data and new tasks over time. The classification emphasizes the adaptation aspect of each technique while acknowledging that FL also serves privacy and resource distribution functions, making it inherently multi-faceted in ECCC systems.
@@
Each learning approach offers unique advantages for different edge-cloud deployment scenarios and data distribution characteristics. Transfer learning excels in scenarios with limited target domain data by leveraging pre-trained models from related domains, achieving faster convergence and reduced computational requirements during training [81]. However, transfer learning performance heavily depends on the similarity between source and target domains, potentially struggling with domain shifts or highly specialized tasks. Federated learning provides exceptional privacy preservation by keeping data distributed across edge devices while enabling collaborative model training, making it ideal for scenarios with sensitive data or regulatory constraints [77]. Nevertheless, FL faces challenges with non-IID data distributions, communication overhead, and convergence difficulties in heterogeneous environments [82]. Continual learning offers superior adaptability to evolving data distributions and new tasks without requiring access to historical data, enabling lifelong learning capabilities crucial for dynamic edge environments [83]. However, CL techniques often struggle with catastrophic forgetting and require sophisticated regularization mechanisms to maintain performance on previous tasks.
@@
The choice of learning paradigm should align with specific data distribution characteristics: (1) Homogeneous data distributions, transfer learning works well when source and target domains share similar characteristics, while federated learning benefits from consistent data patterns across participants; (2) Non-IID data distributions, specialized FL techniques (e.g., FedProx, FedNova) address statistical heterogeneity, while transfer learning may require domain adaptation methods; (3) Temporal data shifts, continual learning excels in adapting to evolving data patterns, while transfer learning may need periodic retraining; (4) Limited local data, transfer learning provides quick adaptation with few samples, while FL can leverage collective knowledge across devices even with sparse local datasets. The optimal selection often involves hybrid approaches that combine multiple paradigms to address specific edge-cloud requirements and constraints. Furthermore, Table III presents a comparative summary of recent research efforts utilizing TL, FL, and CL, highlighting their specific approaches, application domains, and system characteristics within ECCC. Following this, the subsequent sections delve into the transfer learning in Sec. III-B1, federated learning in Sec. III-B2, and continual learning in Sec. III-B3.
@@
TABLE III: Summary of TL, FL, and CL techniques for enhancing AI model adaptability in ECCC systems.
@@
CoLLaRS [84] 2024 FGCS CL AIoT Applications
Li et al. [81] 2024 IEEE TNSM TL Manufacturing
Pan et al. [85] 2024 IEEE Netw. FL Large Model Services
MADDPG [86] 2024 IEEE IoTJ TL 6G Networks
DTS-KD [65] 2024 IEEE IoTJ TL Industrial IoT
Wang et al. [69] 2024 CVPR CL Multimodal Large Language Models
Wang et al. [87] 2024 ESWA FL Rotating Machinery Fault Diagnosis
FedAgg [77] 2024 INFOCOM FL End-Edge-Cloud Systems
OTO [79] 2024 IEEE TPDS TL Cloud-Edge Collaborative DRL
U-Federator [88] 2024 JNCA FL Cloud-Edge-Fog Computing
OTFAC [89] 2024 IEEE IoTJ FL Cloud-Edge Collaborative IoT
FLAT [90] 2024 IEEE ISCC FL Cloud-Edge-Client Continuum
FedITA [91] 2024 AEI FL Industrial Motor Fault Diagnosis
Khuen et al. [92] 2024 IEEE Access FL Precision Aquaculture
Kim et al. [93] 2024 IEEE IoTJ FL Cloud-Edge-Terminal IoT Networks
QuAsyncFL [68] 2024 IEEE IoTJ FL Cloud-Edge-Terminal AIoT
Loconte et al. [94] 2024 FGCS FL Cloud-to-Edge-IoT Continuum
BAFL-DT [82] 2024 IEEE TSC FL Digital Twin Edge-Cloud Continuum
FDRL [95] 2024 IJPR FL Cloud-Edge Manufacturing Systems
Wei et al. [96] 2024 IEEE TPDS FL Edge Clouds
Yang et al. [97] 2024 IEEE IoTJ FL AIoT Data Prediction
FedCAE [98] 2024 IEEE TIE FL Machine Fault Diagnosis
FlexibleFL [99] 2024 INS FL Cloud-Edge Federated Learning
PSPFL [100] 2024 IEEE TMC FL Mobile Edge-Cloud Networks
CCSFLF [101] 2024 CCPE FL Federated Learning Systems
Zhou et al. [16] 2024 ADHOC FL AIoT Quality of Service
CADCL-DTM [102] 2024 CAC CL Rolling Bearings Monitoring
Li et al. [103] 2024 IEEE TMC CL Edge Computing Service Provisioning
Lian et al. [104] 2024 SDI CL ITS Object Detection
EdgeC3 [105] 2024 IEEE TMC CL Distributed Deep Learning
@@ III-B1 Transfer Learning
Pre-trained model adaptation to new tasks or datasets leverages transfer learning principles to reduce training time and data requirements, thereby improving deployment efficiency in resource-constrained environments [81]. For example, in surface defect detection, a model pre-trained on a general image dataset can be fine-tuned with a smaller, specialized dataset to identify specific defects [81]. Similarly, in cloud-device collaborative learning for multimodal large language models (MLLMs), techniques like adapter-based knowledge distillation (AKD) facilitate knowledge transfer from larger cloud models to smaller, compressed device models [69], maintaining performance while minimizing computational burden. In cloud-edge collaborative fault diagnosis, transfer learning addresses data heterogeneity and distribution shifts by fine-tuning source models trained on diverse user data to bridge distribution gaps and enable effective diagnosis on target datasets [87].
@@ III-B2 Model Adaptation in Federated Learning
The explosive growth of FL research has made it an indispensable paradigm in distributed AI systems, driving critical need to explore how model adaptation mechanisms can be effectively implemented within federated frameworks [100]. As one of the most active research domains in distributed machine learning, FL cannot be overlooked when examining model adaptation techniques, particularly given its unique challenges in handling heterogeneous data distributions and diverse edge environments [92]. We focus specifically on model adaptation strategies within FL contexts, exploring how adaptive mechanisms enable models to effectively respond to distributed, non-IID data patterns while maintaining collaborative training capabilities [94]. Recent frameworks demonstrate sophisticated adaptation approaches, such as FedCAE [98] for machine fault diagnosis that adapts models to diverse industrial environments, and QuAsyncFL [68] for AIoT applications that enables adaptive synchronization strategies across heterogeneous devices. Advanced adaptation mechanisms like SFETEC [106] optimize model adaptation for thing-edge-cloud environments by implementing adaptive model splitting strategies that dynamically adjust to varying communication constraints and device capabilities, addressing non-IID data challenges through intelligent partitioning. Security-aware adaptation frameworks such as FlexibleFL [99] integrate adaptive evaluation mechanisms that assess uploaded parameters and participant contributions, enabling models to adapt against evolving poisoning attacks while maintaining collaborative learning effectiveness. Furthermore, adaptive participant selection and topology optimization strategies [96] enhance FL efficiency by dynamically adjusting collaboration patterns, while blockchain integration approaches [82] provide adaptive security mechanisms that evolve with emerging threats, demonstrating how model adaptation principles are fundamental to advancing FL systems in distributed edge-cloud environments.
@@ III-B3 Continual Learning
Incremental adaptation strategies enable models to evolve with dynamic data and emerging tasks while addressing catastrophic forgetting challenges through lifelong learning capabilities [83]. Several strategies, including regularization, replay, and dynamic architectures, mitigate this issue, ensuring robust and adaptable systems. Specifically, continual learning enables digital twins to synchronize with their evolving physical counterparts and sensor data in edge computing, providing accurate predictions [103, 102]. For instance, in intelligent transportation systems (ITS), cloud-edge collaborative continual adaptation frameworks, leveraging visual prompts and knowledge distillation, enhance the generalization of lightweight edge models for object detection in dynamic environments [104]. Similarly, online management frameworks, such as EdgeC3, optimize model aggregation and dynamic offloading for continuous data streams in edge-cloud collaborative continual learning [105]. Cross-edge federated continual learning (Cross-FCL) frameworks, employing parameter decomposition and various cross-edge strategies, address knowledge retention across devices in multiple FL systems, balancing memory and adaptation for new tasks [107].
@@ III-B4 Split Learning
Distributed learning paradigms partition neural network models across multiple computational nodes to enable collaborative training without sharing raw data, addressing computational limitations while preserving privacy [108]. Unlike FL where complete models are trained locally, split learning divides the neural network architecture at specific cut layers, with edge devices typically handling initial feature extraction layers while cloud or edge servers process deeper computational layers [109]. Strategic architectural division effectively addresses the computational limitations of resource-constrained IoT devices while preserving data privacy, as demonstrated in eHealth applications where heterogeneous workload-based resource recommendation models leverage federated split learning (FSL) for smart city deployments [110]. Advanced frameworks like MADRL-based model partitioning systems optimize the collaboration between cloud-edge-device architectures by jointly considering model aggregation control and resource allocation, addressing device heterogeneity through adaptive partitioning strategies rather than uniform approaches [111].
The optimal selection of split points significantly impacts system performance across multiple dimensions including computational load distribution, communication overhead, and privacy preservation. Efficient parallel split learning (EPSL) frameworks demonstrate how intelligent cut layer selection, combined with subchannel allocation and power control optimization, can minimize per-round training latency in resource-constrained wireless edge networks [112]. The SFETEC framework exemplifies optimized split learning for thing-edge-cloud environments, where base models are trained at client devices while core models are processed at edge servers and aggregated at the cloud, significantly reducing communication overhead and training duration compared to traditional federated approaches [106]. Advanced split learning architectures like Split-and-Bridge address class incremental learning challenges by adaptively partitioning neural networks into specialized segments for handling new and old tasks, demonstrating superior performance in knowledge distillation-based continual learning scenarios [113].
Split learning implementations must carefully balance communication efficiency with model performance, particularly in resource-constrained environments. The CSE-FSL framework introduces auxiliary networks for local weight updates while maintaining single server-side models, dramatically reducing gradient transmission requirements and server storage overhead [114]. For wearable IoMT devices, FedSL frameworks enable collaborative healthcare analytics by allowing even low-end devices to participate in model training through strategic model partitioning, while maintaining privacy by avoiding raw data transmission over wireless networks [115]. Comprehensive evaluations comparing FL and split learning in IoT settings reveal that while split learning provides better learning performance under imbalanced data distributions, FL demonstrates superior communication efficiency, highlighting the importance of application-specific optimization strategies [108]. Optimization considerations collectively enable split learning deployment across diverse 6G edge networks, facilitating connected intelligence through distributed model training architectures [109].

### v5

@@ III-A2 Quantization
Precision reduction techniques effectively minimize memory footprint and computational cost by employing lower-precision representations of model parameters [dong2020cdc], making quantized models particularly suitable for resource-constrained devices. Instead of 32-bit floating-point representations, quantization employs lower precision formats such as 8-bit integers, consequently diminishing memory requirements and accelerating computations on these lower-precision numbers. For example, [chen2024implementing] explored model quantization for edge implementation in arrhythmia monitoring to reduce energy consumption. Additionally, QuAsyncFL [liu2024quasyncfl] integrates quantization with asynchronous federated learning to enhance communication efficiency in cloud-edge-terminal collaborations within the AIoT domain.
@@ III-A3 Knowledge Distillation
Knowledge transfer approaches enable compact “student” models to achieve comparable performance with reduced complexity by learning from larger “teacher” models [yao2024gkt], facilitating efficient deployment in resource-constrained environments [tang2024anomaly]. Instead of learning from ground truth labels, the student model learns the teacher model’s output distribution, providing richer information than hard labels and enabling more nuanced representations and better generalization [wang2024clouddevice]. KD benefits ECCC by placing the teacher model in the resource-rich cloud and the distilled student model on edge devices [yang2025growthadaptive], optimizing overall system performance. Agglomerative federated learning (FedAgg) [wu2024agglomerative] uses a bridge sample based online distillation protocol (BSBODP) for knowledge transfer between nodes across different tiers, enabling larger model training while maintaining privacy and flexibility. Guidance-based knowledge transfer (GKT) [yao2024gkt] employs a larger LLM as a “teacher” to generate guidance prompts for a smaller “student” model, enhancing accuracy and speed without fine-tuning. For example, in autonomous driving, a powerful cloud-based LLM can distill its complex decision-making logic into a smaller, faster model suitable for real-time navigation on an in-vehicle computer. Growth-adaptive distillation [yang2025growthadaptive] automatically adjusts the student model structure for efficient, low-power edge deployment. In cloud-edge collaborative reinforcement learning, algorithm-independent knowledge distillation [zeng2024offlinetransferonline] accelerates online RL agent convergence by leveraging pre-trained cloud models and offline agent interactions. A federated bi-directional knowledge distillation strategy [zhang2024cloud] enhances fault detection in manufacturing by integrating quality knowledge from both cloud and edge, strengthening interaction between sub-processes and hierarchical levels.
@@ III-B Model Adaptation
Model adaptation enables AI models to respond to new tasks, datasets, or evolving data distributions in dynamic ECCC environments. Our taxonomy for model adaptation is organized based on the nature of the adaptation challenge being addressed. Accordingly, this leads to a classification into four key paradigms, as depicted in Fig. 5: (1) Transfer learning (Sec. III-B1) focuses on adapting knowledge from a source domain to a target domain, (2) Federated learning (Sec. III-B2) addresses adaptation across distributed, private datasets, (3) Continual learning (Sec. III-B3) handles adaptation to sequential data streams over time, and (4) Split learning (Sec. III-B4) facilitates adaptation by partitioning models across devices. Each paradigm offers a distinct approach to achieving model adaptability under the unique constraints of ECCC.
@@
While some taxonomies treat federated learning as a distinct resource management or privacy technique, our classification positions FL within model adaptation because it fundamentally addresses how models adapt to distributed data patterns and heterogeneous environments, which is a core adaptation challenge in ECCC systems. More specifically, we distinguish three adaptation mechanisms: (1) Domain adaptation techniques (transfer learning) enable models to adapt across different domains or tasks by leveraging pre-existing knowledge; (2) Distributed adaptation methods (federated learning) enable models to adapt collaboratively across multiple data sources while preserving privacy and handling non-IID distributions; (3) Temporal adaptation approaches (continual learning) enable models to adapt incrementally to evolving data and new tasks over time. Notably, this classification emphasizes the adaptation aspect of each technique while acknowledging that FL also serves privacy and resource distribution functions, making it inherently multi-faceted in ECCC systems.
@@
Each learning approach offers unique advantages for different edge-cloud deployment scenarios and data distribution characteristics. Transfer learning excels in scenarios with limited target domain data by leveraging pre-trained models from related domains, achieving faster convergence and reduced computational requirements during training [li2024transfer]. However, transfer learning performance heavily depends on the similarity between source and target domains, potentially struggling with domain shifts or highly specialized tasks. Federated learning provides exceptional privacy preservation by keeping data distributed across edge devices while enabling collaborative model training, making it ideal for scenarios with sensitive data or regulatory constraints [wu2024agglomerative]. Nevertheless, FL faces challenges with non-IID data distributions, communication overhead, and convergence difficulties in heterogeneous environments [qu2024blockchained]. Continual learning offers superior adaptability to evolving data distributions and new tasks without requiring access to historical data, enabling lifelong learning capabilities crucial for dynamic edge environments [yang2024continual]. However, CL techniques often struggle with catastrophic forgetting and require sophisticated regularization mechanisms to maintain performance on previous tasks.
@@
The choice of learning paradigm should align with specific data distribution characteristics: (1) Homogeneous data distributions, transfer learning works well when source and target domains share similar characteristics, while federated learning benefits from consistent data patterns across participants; (2) Non-IID data distributions, specialized FL techniques (e.g., FedProx, FedNova) address statistical heterogeneity, while transfer learning may require domain adaptation methods; (3) Temporal data shifts, continual learning excels in adapting to evolving data patterns, while transfer learning may need periodic retraining; (4) Limited local data, transfer learning provides quick adaptation with few samples, while FL can leverage collective knowledge across devices even with sparse local datasets. Notably, the optimal selection often involves hybrid approaches that combine multiple paradigms to address specific edge-cloud requirements and constraints. Furthermore, Table IV presents a comparative summary of recent research efforts utilizing TL, FL, and CL, highlighting their specific approaches, application domains, and system characteristics within ECCC. Following this, the subsequent sections delve into the transfer learning in Sec. III-B1, federated learning in Sec. III-B2, and continual learning in Sec. III-B3.
@@
TABLE IV: Summary of TL, FL, and CL techniques for enhancing AI model adaptability in ECCC systems.
@@
CoLLaRS [hu2024collars] 2024 FGCS CL AIoT Applications
Li et al. [li2024transfer] 2024 IEEE TNSM TL Manufacturing
Pan et al. [pan2024cloudedge] 2024 IEEE Netw. FL Large Model Services
MADDPG [she2024efficient] 2024 IEEE IoTJ TL 6G Networks
DTS-KD [tang2024anomaly] 2024 IEEE IoTJ TL Industrial IoT
Wang et al. [wang2024clouddevice] 2024 CVPR CL Multimodal Large Language Models
Wang et al. [wang2024cloudedgea] 2024 ESWA FL Rotating Machinery Fault Diagnosis
FedAgg [wu2024agglomerative] 2024 INFOCOM FL End-Edge-Cloud Systems
OTO [zeng2024offlinetransferonline] 2024 IEEE TPDS TL Cloud-Edge Collaborative DRL
U-Federator [ali2024universal] 2024 JNCA FL Cloud-Edge-Fog Computing
OTFAC [gan2024optimal] 2024 IEEE IoTJ FL Cloud-Edge Collaborative IoT
FLAT [garofalo2024webcentric] 2024 IEEE ISCC FL Cloud-Edge-Client Continuum
FedITA [he2024fedita] 2024 AEI FL Industrial Motor Fault Diagnosis
Khuen et al. [khuencheng2024integration] 2024 IEEE Access FL Precision Aquaculture
Kim et al. [kim2024collaborative] 2024 IEEE IoTJ FL Cloud-Edge-Terminal IoT Networks
QuAsyncFL [liu2024quasyncfl] 2024 IEEE IoTJ FL Cloud-Edge-Terminal AIoT
Loconte et al. [loconte2024expanding] 2024 FGCS FL Cloud-to-Edge-IoT Continuum
BAFL-DT [qu2024blockchained] 2024 IEEE TSC FL Digital Twin Edge-Cloud Continuum
FDRL [wang2024federated] 2024 IJPR FL Cloud-Edge Manufacturing Systems
Wei et al. [wei2024joint] 2024 IEEE TPDS FL Edge Clouds
Yang et al. [yang2024differentially] 2024 IEEE IoTJ FL AIoT Data Prediction
FedCAE [yu2024fedcae] 2024 IEEE TIE FL Machine Fault Diagnosis
FlexibleFL [zhao2024flexiblefl] 2024 INS FL Cloud-Edge Federated Learning
PSPFL [zhou2024accelerating] 2024 IEEE TMC FL Mobile Edge-Cloud Networks
CCSFLF [zhou2024ccsflf] 2024 CCPE FL Federated Learning Systems
Zhou et al. [zhou2024enhancing] 2024 ADHOC FL AIoT Quality of Service
CADCL-DTM [li2024conditionadaptive] 2024 CAC CL Rolling Bearings Monitoring
Li et al. [li2024digital] 2024 IEEE TMC CL Edge Computing Service Provisioning
Lian et al. [lian2024cloudedge] 2024 SDI CL ITS Object Detection
EdgeC3 [lin2024online] 2024 IEEE TMC CL Distributed Deep Learning
@@ III-B1 Transfer Learning
Pre-trained model adaptation to new tasks or datasets leverages transfer learning principles to reduce training time and data requirements, thereby improving deployment efficiency in resource-constrained environments [li2024transfer]. For example, in surface defect detection, a model pre-trained on a general image dataset can be fine-tuned with a smaller, specialized dataset to identify specific defects [li2024transfer]. Similarly, in cloud-device collaborative learning for multimodal large language models (MLLMs), techniques like adapter-based knowledge distillation (AKD) facilitate knowledge transfer from larger cloud models to smaller, compressed device models [wang2024clouddevice], maintaining performance while minimizing computational burden. In cloud-edge collaborative fault diagnosis, transfer learning addresses data heterogeneity and distribution shifts by fine-tuning source models trained on diverse user data to bridge distribution gaps and enable effective diagnosis on target datasets [wang2024cloudedgea].
@@ III-B2 Model Adaptation in Federated Learning
The explosive growth of FL research has made it an indispensable paradigm in distributed AI systems, driving critical need to explore how model adaptation mechanisms can be effectively implemented within federated frameworks [zhou2024accelerating]. As one of the most active research domains in distributed machine learning, FL cannot be overlooked when examining model adaptation techniques, particularly given its unique challenges in handling heterogeneous data distributions and diverse edge environments [khuencheng2024integration]. We focus specifically on model adaptation strategies within FL contexts, exploring how adaptive mechanisms enable models to effectively respond to distributed, non-IID data patterns while maintaining collaborative training capabilities [loconte2024expanding]. Recent frameworks demonstrate sophisticated adaptation approaches, such as FedCAE [yu2024fedcae] for machine fault diagnosis that adapts models to diverse industrial environments, and QuAsyncFL [liu2024quasyncfl] for AIoT applications that enables adaptive synchronization strategies across heterogeneous devices. Advanced adaptation mechanisms like SFETEC [le2024sfetec] optimize model adaptation for thing-edge-cloud environments by implementing adaptive model splitting strategies that dynamically adjust to varying communication constraints and device capabilities, addressing non-IID data challenges through intelligent partitioning. Security-aware adaptation frameworks such as FlexibleFL [zhao2024flexiblefl] integrate adaptive evaluation mechanisms that assess uploaded parameters and participant contributions, enabling models to adapt against evolving poisoning attacks while maintaining collaborative learning effectiveness. Despite its privacy-preserving architecture, FL remains vulnerable to model privacy attacks such as membership inference, which are discussed in detail in LABEL:sec:supp-model-privacy-attacks. Furthermore, adaptive participant selection and topology optimization strategies [wei2024joint] enhance FL efficiency by dynamically adjusting collaboration patterns, while blockchain integration approaches [qu2024blockchained] provide adaptive security mechanisms that evolve with emerging threats, demonstrating how model adaptation principles are fundamental to advancing FL systems in distributed edge-cloud environments.
@@ III-B3 Continual Learning
Incremental adaptation strategies enable models to evolve with dynamic data and emerging tasks while addressing catastrophic forgetting challenges through lifelong learning capabilities [yang2024continual]. For example, in a smart factory, a continual learning model for predictive maintenance can adapt to new machine behaviors and fault patterns over time without being completely retrained, ensuring the system remains accurate as equipment ages. Several strategies, including regularization, replay, and dynamic architectures, mitigate this issue, ensuring robust and adaptable systems. Specifically, continual learning enables digital twins to synchronize with their evolving physical counterparts and sensor data in edge computing, providing accurate predictions [li2024digital, li2024conditionadaptive]. For instance, in intelligent transportation systems (ITS), cloud-edge collaborative continual adaptation frameworks, leveraging visual prompts and knowledge distillation, enhance the generalization of lightweight edge models for object detection in dynamic environments [lian2024cloudedge]. Similarly, online management frameworks, such as EdgeC3, optimize model aggregation and dynamic offloading for continuous data streams in edge-cloud collaborative continual learning [lin2024online]. Cross-edge federated continual learning (Cross-FCL) frameworks, employing parameter decomposition and various cross-edge strategies, address knowledge retention across devices in multiple FL systems, balancing memory and adaptation for new tasks [zhang2024crossfcl].
@@ III-B4 Split Learning
Distributed learning paradigms partition neural network models across multiple computational nodes to enable collaborative training without sharing raw data, addressing computational limitations while preserving privacy where early layers typically run on edge devices for low-latency feature extraction and deeper layers are offloaded to servers [gao2020endend, lin2024split]. Strategic architectural division effectively addresses device limitations and privacy, as seen in federated split learning (FSL) for smart cities [ahmed2024heterogeneous] and MADRL-based partitioning systems that optimize cloud-edge collaboration [fan2025madrlbased]. Furthermore, selecting optimal split points minimizes latency and communication overhead, with frameworks like EPSL [lin2024efficient] and SFETEC [le2024sfetec] demonstrating efficiency in wireless and thing-edge-cloud environments respectively. Moreover, specialized architectures address incremental learning challenges [kim2021splitbridge], while implementations like CSE-FSL [mu2025federated] and FedSL [ni2024fedsl] balance communication efficiency with model performance, enabling secure participation of low-end devices in diverse 6G networks [lin2024split].